# Graph-Theoretic Analysis of $n$-Replica Time Evolution in the Brownian Gaussian Unitary Ensemble

### v1

@@ 5 General Cases
$X^{(\boldsymbol{\alpha}_{j})}$ is a set of 2-swap operators submitting to the condition $\boldsymbol{\alpha}_{j}$. Our primary goal is to deal with the following graph expression
$\sum_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})c\left[\prod_{j=1}^{m}(X+\bar{X})^{(\boldsymbol{\alpha}_{j})}\right]\ ,$ (129)
and assign them into appropriate categories.
@@
In practice, we can always find a permutation $\sigma\in S_{n}$ as a product of cycles $\{\sigma_{i}\}_{i=1,2,\ldots,g}$ with length $1<l_{1}\leq l_{2}\ldots\leq l_{g}$ such that:
@@
Then for a certain $p$, we first analyze all permutation categories appearing in the products $\prod_{j=1}^{m}X^{(\alpha_{j})},m=0,1,2,\ldots$, which gives us a division of the permutation group $S_{n}$
@@
Finally when $p\geq 2$, we need to consider removing the permutation redundancy of the conditions and the gauge symmetry of the graph representation. In other words, we need to module some equivalent relations.
@@ 5.1 Example: $n=4$
Here we take $n=4$ as an example. For different number of connected pairs, the number of graphs should be $N_{4,p}={\{576,9216,20736,9216,576\}}$. We then define
@@
For $p=0$, there is no restriction on $X$, so we have the division of $S_{4}$ as
@@
Since there is no permutation redundancy or gauge redundancy, we have ${N_{X}(N_{X}+1)\over 2}=15$ independent graph categories:
@@
$S_{4}=I+X^{(0)}+X^{(1)}+X_{[3]}^{(0)}+X_{[3]}^{(1)}+X_{[2,2]}+X_{[4]}\equiv\cup_{\alpha=1}^{7}X_{\alpha}\ .$ (137)
Since $p<2$, there is no permutation or gauge redundancy, so we have ${7\times 8\over 2}=28$ categories:
$F_{1,X_{\alpha}\bar{X}_{\beta}}\equiv\sum_{u}\sum_{x\in X_{\alpha}\bar{X}_{\beta}\cup X_{\beta}\bar{X}_{\alpha}}c_{u}c[x].$ (138)
@@
$X=X^{(00)}+X^{(01)}+X^{(10)}+X^{(11)},~{}X_{[3]}^{(00)}=\emptyset,~{}X_{[3]}=X_{[3]}^{(01)}+X^{(10)}_{[3]}+X_{[3]}^{(11)}\ ,$ (139)
so we obtain the division to be
@@
Naively, we have ${10\times 11\over 2}=55$ categories. But here we need to remove the permutation redundancy. For an example, changing the order of conditions $\alpha_{j}$, we have
$\mathsf{F}_{2,X}^{(01)}=\mathsf{F}_{2,X}^{(10)},~{}\mathsf{F}_{2,X\bar{X}}^{(00,01)}=\mathsf{F}_{2,X\bar{X}}^{(00,10)}\ .$ (141)
@@
As for the gauge redundancy, here we impose the equivalence $X^{(11)}\bar{X}^{(11)}\sim I$, so immediately we have
@@
Naively, there are ${12\times 13\over 2}=96$ categories. The permutation redundancy is similar to $p=2$ case and the gauge group is generated by
@@
which gives the division as
@@
Notice the gauge group is generated by each permutation like $X_{e}\bar{X}_{e}$. So we can fix the order of $\bar{i}$ to be $\bar{1}\bar{2}\bar{3}\bar{4}$, and we just need to consider the categories
@@
Notice the permutation redundancy, we have
@@
so we only have five categories left
@@
6 Discussions and Outlook
In this paper we analyze the time evolution of the $n$-replica physics of the BGUE model in a graph-based framework. This approach proves to be particularly effective for lower values of $n$, as explicitly demonstrated in the $n=2$ and $n=3$ cases. We also provide a systematic strategy that extends to generic $n$. Our construction highlights how the increasing complexity of graph structures can be managed by organizing them into well-defined categories, which will simplify the calculations of two-point functions, OTOC, and other relevant observables.
The general formulation outlined in Sec. 5 suggests that this method can, in principle, be extended to arbitrary $n$. However, as $n$ increases, the growing number of graph structures presents significant challenges. Addressing these challenges will require the development of more efficient techniques for classifying, encoding, and manipulating graphs at large $n$. An important direction for future research is to explore algorithmic approaches to systematically handle this complexity. Moreover, further investigation into the implications of this framework for Brownian disordered systems and quantum information theory may uncover deeper structural insights into the role of replica dynamics in random matrix models.
A key insight is that the factor $w$, which encodes the dependence on the system spectrum, commutes with the graph-based operators. This allows us to analyze the system with infinite-temperature GUE noise without introducing additional complexity.
It is worth noting that the techniques developed in this paper can also be applied to other models, such as the Brownian Gaussian Orthogonal Ensemble (BGOE) and the Brownian Gaussian Symplectic Ensemble (BGSE). In particular, for BGOE, we have:
@@ 6 Discussions and Outlook
We find there are new operators such as $P_{ij},P_{\bar{i}\bar{j}},X_{i\bar{j}}$. Hence an individual graph appearing in $\mathcal{L}_{n}^{r}$ can be written as
@@
As a result, when $n$ increases, the introduction of new operators in BGOE and BGSE further complicates the graph structures, leading to a greater number of individual graphs that become increasingly challenging to systematically organize them into appropriate categories. While this classification seems to be complicated, it is, at its core, a well-structured problem. As discussed in the introduction, the method presented in this paper is fundamentally equivalent to the approach outlined in Guo:2024zmr . Moreover, the inherent symmetry of $\mathcal{L}_{n}$ in the BGUE, BGOE, and BGSE cases simplifies its diagonalization, which is closely related to a well-studied mathematical problem: the decomposition of the direct product of fundamental representations. Given this, we expect that the techniques developed in this work can also be successfully extended to models without obvious symmetry, further demonstrating the power and versatility of the proposed framework.
@@ Acknowledgment
The authors would like to thank Haifeng Tang, Yingyu Yang, and Yanyuan Li for their valuable discussions and comments during the preparation of this work. CP, JY and TL are supported by NSFC NO. 12175237, CP is also supported in part by NSFC NO. 12447108, NSFC NO. 12247103, the Fundamental Research Funds for the Central Universities, and funds from the Chinese Academy of Sciences.
@@ Appendix A Expression of $\mathcal{U}_{3}(t)$
As discussed in Sec.4, we write $\mathcal{U}_{3}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}$. Here we list the explicit expressions for $f_{a}$. For simplicity, we denote
$f_{a}(t)=\boldsymbol{v}_{a}\cdot\boldsymbol{\xi}(t)\ ,$ (155)
@@
$\boldsymbol{\xi}(t)=\left\{e^{-Jt},e^{-2Jt},e^{-\frac{2(D+1)Jt}{D}},e^{-\frac{2(D-1)Jt}{D}},e^{-3Jt},e^{-\frac{3(D+1)Jt}{D}},e^{-\frac{3(D-1)Jt}{D}},e^{-\frac{3(D+2)Jt}{D}},e^{-\frac{3(D-2)Jt}{D}},1\right\}\ ,$ (156)
@@
$\boldsymbol{v}_{1}$ =$=\left\{0,0,0,0,\frac{1}{2},\frac{2}{9},\frac{2}{9},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{2}$ =$=\left\{0,0,0,0,0,\frac{1}{9},-\frac{1}{9},\frac{1}{36},-\frac{1}{36},0\right\},$
$\boldsymbol{v}_{3}$ =$=\left\{0,0,0,0,-\frac{1}{18},0,0,\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{4}$ =$=\left\{0,0,0,0,-\frac{1}{6},\frac{1}{18},\frac{1}{18},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{5}$ =$=\left\{0,0,0,0,0,-\frac{1}{18},\frac{1}{18},\frac{1}{36},-\frac{1}{36},0\right\},$
$\boldsymbol{v}_{6}$ =$=\left\{0,0,0,0,\frac{1}{6},-\frac{1}{9},-\frac{1}{9},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{7}$ =$=\Bigg{\{}0,\frac{5-D^{2}}{18D-2D^{3}},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},$
@@
$\boldsymbol{v}_{8}$ =$=\Bigg{\{}0,0,\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{-D^{2}+3D+2}{4\left(D^{3}-3D^{2}-6D+8\right)},\frac{4}{36-9D^{2}},$
@@
$\boldsymbol{v}_{9}$ =$=\Bigg{\{}0,\frac{1}{18-2D^{2}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},$
@@
$\boldsymbol{v}_{10}$ =$=\Bigg{\{}0,-\frac{D^{2}-5}{2D\left(D^{2}-9\right)},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},\frac{20-D^{2}}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{11}$ =$=\Bigg{\{}0,\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},-\frac{D^{2}+4}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{12}$ =$=\Bigg{\{}0,\frac{1}{D\left(D^{2}-9\right)},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{8-D^{2}}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{13}$ =$=\Bigg{\{}0,-\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{4-3D^{2}}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{14}$ =$=\Bigg{\{}0,\frac{1}{2\left(D^{2}-9\right)},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{2}{9\left(D^{2}-4\right)},$
@@
$\boldsymbol{v}_{15}$ =$=\Bigg{\{}0,0,\frac{1}{2(D-2)(D+1)(D+4)},-\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{1}{36-9D^{2}},\frac{1}{18D+18},$
@@
$\boldsymbol{v}_{16}$ =$=\Bigg{\{}0,\frac{1}{9D-D^{3}},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},-\frac{8-3D^{2}}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{17}$ =$=\Bigg{\{}\frac{D^{4}-8D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3-D^{2}}{D^{2}\left(D^{2}-9\right)},-\frac{D(D+3)}{2\left(D^{4}+5D^{3}-20D-16\right)},-\frac{(D-3)D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
@@
$\boldsymbol{v}_{18}$ =$=\Bigg{\{}\frac{1}{9D-D^{3}},-\frac{1}{18D-2D^{3}},-\frac{D}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D}{4\left(D^{3}-3D^{2}-6D+8\right)},-\frac{2}{36D-9D^{3}},$
@@
$\boldsymbol{v}_{19}$ =$=\Bigg{\{}\frac{1}{9D-D^{3}},\frac{2}{D\left(D^{2}-9\right)},\frac{1}{(D-2)(D+1)(D+4)},\frac{1}{D^{3}-3D^{2}-6D+8},\frac{4}{36D-9D^{3}},$
@@
$\boldsymbol{v}_{20}$ =$=\Bigg{\{}\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{D^{2}-3}{2D^{2}\left(D^{2}-9\right)},\frac{D^{2}+3D+4}{-4D^{4}-20D^{3}+80D+64},$
@@
$\boldsymbol{v}_{21}$ =$=\Bigg{\{}\frac{D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3}{D^{2}\left(D^{2}-9\right)},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{-2D^{4}+10D^{3}-40D+32},$
@@
$\boldsymbol{v}_{22}$ =$=\Bigg{\{}\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{3}{18D^{2}-2D^{4}},\frac{3D+4}{4\left(D^{4}+5D^{3}-20D-16\right)},\frac{4-3D}{4\left(D^{4}-5D^{3}+20D-16\right)},$
@@
$\boldsymbol{v}_{23}$ =$=\Bigg{\{}-\frac{5}{D^{5}-13D^{3}+36D},\frac{1}{9D-D^{3}},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
@@
$\boldsymbol{v}_{24}$ =$=\Bigg{\{}\frac{3}{9D-D^{3}},0,\frac{3}{2(D-2)(D+1)(D+4)},\frac{3}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{2}{12D-3D^{3}},0,0,$
@@
$\boldsymbol{v}_{25}$ =$=\Bigg{\{}\frac{5}{D^{4}-13D^{2}+36},0,\frac{1}{2\left(D^{3}+7D^{2}+14D+8\right)},\frac{1}{-2D^{3}+14D^{2}-28D+16},0,0,0,$
@@
$\boldsymbol{v}_{26}$ =$=\Bigg{\{}-\frac{15}{D^{5}-13D^{3}+36D},0,-\frac{3}{D^{4}+5D^{3}-20D-16},\frac{3}{D^{4}-5D^{3}+20D-16},-\frac{1}{12D-3D^{3}},$

### v2

@@ 5 General case
In practice, we can always write a permutation $\sigma\in S_{n}$ as a product of cycles $\{\sigma_{i}\}_{i=1,2,\ldots,g}$ with length $1<l_{1}\leq l_{2}\ldots\leq l_{g}$:
@@
Then for certain $p$, we first analyze all permutation categories appearing in the products $\prod_{j=1}^{m}X^{(\alpha_{j})},m=0,1,2,\ldots$, which gives a division of the permutation group $S_{n}$
@@
Finally when $p\geq 2$, we need to consider removing the permutation redundancy of the conditions and the gauge symmetry of the graph representation, i.e, we need to module some equivalence relation.
@@ 5.1 Example: $n=4$
Here we take $n=4$ as an example, then $N_{4,p}={\{576,9216,20736,9216,576\}}$. We then define
@@
Due to $p=0$, there is no constricts on $X$, so we have the division of $S_{4}$
@@
Since there is no gauge redundancy and condition permutation redundancy, we have ${N_{X}(N_{X}+1)\over 2}=15$ graph categories:
@@
$S_{4}=I+X^{(0)}+X^{(1)}+X_{[3]}^{(0)}+X_{[3]}^{(1)}+X_{[2,2]}+X_{[4]}\equiv\cup_{\alpha}X_{\alpha}\ .$ (131)
Due to $p<2$, there is no redundancy, so we have ${7\times 8\over 2}=28$ categories: $F_{1,X_{\alpha}\bar{X}_{\beta}}\equiv\sum_{u,x\in X_{\alpha}\bar{X}_{\beta}}c_{u}c[x]$.
@@
$X=X^{(00)}+X^{(01)}+X^{(10)}+X^{(11)},X_{[3]}^{(00)}=\emptyset,X_{[3]}=X_{[3]}^{(01)}+X^{(10)}_{[3]}+X_{[3]}^{(11)}\ ,$ (132)
so we obtain the division
@@
Naively, we have ${10\times 11\over 2}=55$ categories. But here we need to remove the condition redundancy and gauge redundancy. For an example, changing the order of conditions, we have
$\mathsf{F}_{2,X}^{(01)}=\mathsf{F}_{2,X}^{(10)},\mathsf{F}_{2,X\bar{X}}^{(00,01)}=\mathsf{F}_{2,X\bar{X}}^{(00,10)}\ .$ (134)
@@
As for the gauge redundancy, here we impose the equivalence $X^{(11)}\bar{X}^{(11)}\sim I$, so directly we have
@@
Naively, there are ${12\times 13\over 2}=96$ categories. The condition permutation redundancy is similar to $p=2$ case and the gauge group is generated by
@@
which gives the division
@@
Notice the gauge group is generated by any permutation like $X_{e}\bar{X}_{e},\forall e$. So we can fix the order of $\bar{i}$ to be $\bar{1}\bar{2}\bar{3}\bar{4}$, then we just need to consider the categories
@@
Notice the permutation redundancy of the condition, we have
@@
so we only have five categories
@@
6 Discussion and Outlook
The graph-based framework developed in this work provides a powerful approach for analyzing the $n$-replica physics of the BGUE model. The explicit constructions for the $n=2$ and $n=3$ cases demonstrate how the increasing complexity of the graphs can be managed by organizing them into distinct categories. The general formulation outlined in Section 5 suggests that this approach can be extended to arbitrary $n$, though the proliferation of graph structures poses a significant challenge. Developing efficient methods for packing and manipulating the graphs will be an important direction for future research.
A key insight is that the factor $w$, which encodes the dependence on the system spectrum, commutes with the graph-based operators. This allows for the analysis of the system with infinite-temperature GUE noise without introducing additional complexity.
The techniques discussed in this paper can be applied to other models, such as the Brownian Gaussian Orthogonal Ensemble (BGOE) and Brownian Gaussian Symplectic Ensemble (BGSE). For BGOE, we have
@@ 6 Discussion and Outlook
We find there are new operators $P_{ij},P_{\bar{i}\bar{j}},X_{i\bar{j}}$, so an individual graph appearing in $\mathcal{L}_{n}^{r}$ can be written as
@@
So there will be more individual graphs, which can be challenging to organize into appropriate categories. While this task may seem intricate, it is fundamentally a straightforward endeavor. However, as discussed in the introduction, the method outlined in this paper is essentially equivalent to the approach detailed in Guo:2024zmr . The inherent symmetry of $\mathcal{L}_{n}$ for BGUE, BGOE and BGSE simplifies its diagonalization, a well-studied problem in the field of mathematics: the decomposition of the direct product of the fundamental representations. We therefore anticipate that the techniques employed in this paper can be successfully applied to models lacking obvious symmetry, thereby highlighting the power and versatility of the proposed method.
@@ Acknowledgment
The author would like to thank Haifeng Tang, Cheng Peng, Yingyu Yang, and Yanyuan Li for their valuable discussions and comments during the preparation of this work.
@@ Appendix A Expression of $\mathcal{U}_{3}(t)$
As discussed in Sec.4, we write $\mathcal{U}_{3}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}$. Here we list the explicit expressions for $f_{a}$’s. For simplicity, we denote
$f_{a}(t)=\boldsymbol{v}_{a}\cdot\boldsymbol{\xi}(t)$ (148)
@@
$\boldsymbol{\xi}(t)=\left\{e^{-Jt},e^{-2Jt},e^{-\frac{2(D+1)Jt}{D}},e^{-\frac{2(D-1)Jt}{D}},e^{-3Jt},e^{-\frac{3(D+1)Jt}{D}},e^{-\frac{3(D-1)Jt}{D}},e^{-\frac{3(D+2)Jt}{D}},e^{-\frac{3(D-2)Jt}{D}},1\right\}$ (149)
@@
$\boldsymbol{v}_{1}$ $=\left\{0,0,0,0,\frac{1}{2},\frac{2}{9},\frac{2}{9},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{2}$ $=\left\{0,0,0,0,0,\frac{1}{9},-\frac{1}{9},\frac{1}{36},-\frac{1}{36},0\right\},$
$\boldsymbol{v}_{3}$ $=\left\{0,0,0,0,-\frac{1}{18},0,0,\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{4}$ $=\left\{0,0,0,0,-\frac{1}{6},\frac{1}{18},\frac{1}{18},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{5}$ $=\left\{0,0,0,0,0,-\frac{1}{18},\frac{1}{18},\frac{1}{36},-\frac{1}{36},0\right\},$
$\boldsymbol{v}_{6}$ $=\left\{0,0,0,0,\frac{1}{6},-\frac{1}{9},-\frac{1}{9},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{7}$ $=\Bigg\{0,\frac{5-D^{2}}{18D-2D^{3}},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},$
@@
$\boldsymbol{v}_{8}$ $=\Bigg\{0,0,\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{-D^{2}+3D+2}{4\left(D^{3}-3D^{2}-6D+8\right)},\frac{4}{36-9D^{2}},$
@@
$\boldsymbol{v}_{9}$ $=\Bigg\{0,\frac{1}{18-2D^{2}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},$
@@
$\boldsymbol{v}_{10}$ $=\Bigg\{0,-\frac{D^{2}-5}{2D\left(D^{2}-9\right)},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},\frac{20-D^{2}}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{11}$ $=\Bigg\{0,\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},-\frac{D^{2}+4}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{12}$ $=\Bigg\{0,\frac{1}{D\left(D^{2}-9\right)},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{8-D^{2}}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{13}$ $=\Bigg\{0,-\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{4-3D^{2}}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{14}$ $=\Bigg\{0,\frac{1}{2\left(D^{2}-9\right)},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{2}{9\left(D^{2}-4\right)},$
@@
$\boldsymbol{v}_{15}$ $=\Bigg\{0,0,\frac{1}{2(D-2)(D+1)(D+4)},-\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{1}{36-9D^{2}},\frac{1}{18D+18},$
@@
$\boldsymbol{v}_{16}$ $=\Bigg\{0,\frac{1}{9D-D^{3}},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},-\frac{8-3D^{2}}{72D-18D^{3}},$
@@
$\boldsymbol{v}_{17}$ $=\Bigg\{\frac{D^{4}-8D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3-D^{2}}{D^{2}\left(D^{2}-9\right)},-\frac{D(D+3)}{2\left(D^{4}+5D^{3}-20D-16\right)},-\frac{(D-3)D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
@@
$\boldsymbol{v}_{18}$ $=\Bigg\{\frac{1}{9D-D^{3}},-\frac{1}{18D-2D^{3}},-\frac{D}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D}{4\left(D^{3}-3D^{2}-6D+8\right)},-\frac{2}{36D-9D^{3}},$
@@
$\boldsymbol{v}_{19}$ $=\Bigg\{\frac{1}{9D-D^{3}},\frac{2}{D\left(D^{2}-9\right)},\frac{1}{(D-2)(D+1)(D+4)},\frac{1}{D^{3}-3D^{2}-6D+8},\frac{4}{36D-9D^{3}},$
@@
$\boldsymbol{v}_{20}$ $=\Bigg\{\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{D^{2}-3}{2D^{2}\left(D^{2}-9\right)},\frac{D^{2}+3D+4}{-4D^{4}-20D^{3}+80D+64},$
@@
$\boldsymbol{v}_{21}$ $=\Bigg\{\frac{D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3}{D^{2}\left(D^{2}-9\right)},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{-2D^{4}+10D^{3}-40D+32},$
@@
$\boldsymbol{v}_{22}$ $=\Bigg\{\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{3}{18D^{2}-2D^{4}},\frac{3D+4}{4\left(D^{4}+5D^{3}-20D-16\right)},\frac{4-3D}{4\left(D^{4}-5D^{3}+20D-16\right)},$
@@
$\boldsymbol{v}_{23}$ $=\Bigg\{-\frac{5}{D^{5}-13D^{3}+36D},\frac{1}{9D-D^{3}},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
@@
$\boldsymbol{v}_{24}$ $=\Bigg\{\frac{3}{9D-D^{3}},0,\frac{3}{2(D-2)(D+1)(D+4)},\frac{3}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{2}{12D-3D^{3}},0,0,$
@@
$\boldsymbol{v}_{25}$ $=\Bigg\{\frac{5}{D^{4}-13D^{2}+36},0,\frac{1}{2\left(D^{3}+7D^{2}+14D+8\right)},\frac{1}{-2D^{3}+14D^{2}-28D+16},0,0,0,$
@@
$\boldsymbol{v}_{26}$ $=\Bigg\{-\frac{15}{D^{5}-13D^{3}+36D},0,-\frac{3}{D^{4}+5D^{3}-20D-16},\frac{3}{D^{4}-5D^{3}+20D-16},-\frac{1}{12D-3D^{3}},$